OpenStudy (anonymous):

Today only, a table is being sold for $286 . This is 67% of its regular price. What was the price yesterday?

OpenStudy (anonymous):

$477.62

OpenStudy (anonymous):

286(67)=19162 19162/100=191.62 286+191.62=477.62

OpenStudy (anonymous):

wrong solution mmbuckaroos....

OpenStudy (anonymous):

u added 67% of 286 into it...but we have to find that what is the number whose 67% is 286...the solution is... \[y*67/100=286\]

OpenStudy (anonymous):

Yep sorry I miss read the question

OpenStudy (anonymous):

and the answer is 426...approx..